OpenStudy (nathanjhw):

Use separation of variables to solve the differential equation dy/dx = (1+x)/xy with initial condition y(1)=-2.

OpenStudy (nathanjhw):

1/2 y^2 = ln lxl +x + 1 1/2 y^2 = 1/2(x+1)^2 1/2 y^2 = (x+1)^2 +3 1/2 y^2 = ln lxl +x -3 This one can be separated, but it's impossible to solve using techniques of antidifferentiation.

OpenStudy (nathanjhw):

@freckles

OpenStudy (nathanjhw):

@zimmah

OpenStudy (nathanjhw):

I believe it is either the first one or the fourth one.

OpenStudy (nathanjhw):

Actually I figured it out, it is the first one.

OpenStudy (anonymous):

let me help

OpenStudy (nathanjhw):

I found out the answer. It's the first one.

OpenStudy (anonymous):

wait lets prove it

OpenStudy (anonymous):

dy/dx = (1+x)/xy you were told to solve

OpenStudy (anonymous):

\[ dy/dx = (1+x)/xy \]

OpenStudy (anonymous):

first try to seprate all x to one side and all y to the other

OpenStudy (anonymous):

\[ydy=(1+x)dx/x \]

OpenStudy (anonymous):

now intigrate

OpenStudy (anonymous):

\[\int\limits_{}^{}ydy=\int\limits_{}^{}(1+x)/x dx\]

OpenStudy (anonymous):

but recall that

OpenStudy (anonymous):

\[(1+x)/x=1/x+x/x \]

OpenStudy (anonymous):

\[y^2/2=lnx+x +c\]

OpenStudy (anonymous):

guess u don't need help